Thomas drove 60 kilometers at a constant rate of speed. If he had driven 40 kph faster, he would have traveled the same distance for 24 minutes less. How fast did he drive?
Answer by checkley71(8403)

60/x=60/(x+40)+.4
60/x=(60+.4x+16)/(x+40) now cross multiply
60(x+40)=x(60+.4x+16)
60x+2400=60x+.4x^2+16x
.4x^2+16x-2400=0
x^2+40x-6000=0
(x-60)(x+100)=0
x-60=0
x=60 kph. is the speed.
PROOF:
60/60=60/(60+40)+.4
1=(60+40)/100
1=1